# Intent Prediction-Driven Model Predictive Control for UAV Planning and Navigation in Dynamic Environments

### v1

@@ 2 $N_{\text{pred}}\leftarrow\text{Prediction Time Steps}$;
$p^{o},v^{o},s^{o}\leftarrow\textbf{obsInfo}(o)$ ;
  $\triangleright$ pos, vel, size
@@
13                     $\mathcal{S}_{\text{control}}\leftarrow{\textbf{getLinearControlSet}}(v^{o})$
@@
15                          $\mathcal{S}_{\text{control}}\leftarrow{\textbf{getTurningControlSet}}(v^{o},\mathbb{I}^{o})$
@@
III-C MPC Formulation for Trajectory Generation
@@
The collision constraints (Eqn. III-C) ensure that the robot avoids collisions with static and dynamic obstacles. Static obstacles are represented in an occupancy voxel map, while dynamic obstacles are detected as axis-aligned 3D bounding boxes. To simplify the expression of the collision constraints, we unify the representations by converting static obstacles into oriented 3D bounding boxes with a proposed hierarchical clustering method. Initially, surrounding occupancy voxels are collected as a point cloud, and the DBSCAN clustering algorithm is applied as the first level to determine the rough centroid and dimensions of the static obstacles. This step ensures that all static obstacle voxels are enclosed within one of the DBSCAN-clustered bounding boxes, though these boxes may include empty spaces. To refine these DBSCAN-clustered bounding boxes, we then iteratively apply the K-means++ algorithm, starting with two initial centroids at the corners of the bounding boxes and continuing until the point-cloud density within the boxes exceeds a specified threshold. Finally, the 2D orientation angle $\phi$ of each static obstacle bounding box is adjusted in a discretized manner to further maximize point-cloud density. With both static and dynamic obstacles uniformly represented as 3D bounding boxes, the collision constraint for each obstacle can be expressed as:
@@
III-D Intent-Based Trajectory Planning
@@ 18           $\textbf{return}\ \sigma_{\text{output}}$;
With the dynamic obstacle trajectory prediction and MPC formulation introduced, this section presents the intent-based trajectory planning algorithm. The proposed algorithm (Alg. 2) generates multiple trajectories corresponding to different obstacle intent combinations and selects the optimal one for execution based on an evaluation system. Initially, static obstacles are clustered into a set of bounding boxes $\mathcal{S}_{so}$, and the set of risky dynamic obstacles $\mathcal{S}_{do}$ near the robot is identified (Lines 2-2). Next, the top $N_{ic}$ intent combinations $S_{ic}$ for different obstacles are determined based on their current intent probability distributions. Each intent combination, denoted as $\mathcal{IC}$ in Line 2, represents one possible future action scenario for all obstacles, with the algorithm focusing on the most likely $N_{ic}$ scenarios. For a single intent combination, the trajectory-level prediction is obtained for each obstacle (Line 2), and MPC is used to generate a collision-free trajectory $\sigma_{\text{safeTraj}}$ for navigation (Line 2). Finally, the score of the generated trajectory is calculated by combining the intent combination probability $P_{ic}$ with the raw score from the evaluation system (Lines 2-2). The algorithm selects the highest-score trajectory for execution.
@@
Both the consistency score and the detouring score are capped at a maximum allowable value to prevent extremely high scores when the denominators approach zero. The safety score is determined by the distance to static and dynamic obstacles under the most likely intent combination:
@@
IV Result and Discussion
To evaluate the proposed method, we conduct simulation experiments and physical flight tests in various dynamic environments. The system is implemented in C++ and ROS and runs on an NVIDIA Orin NX onboard computer from our quadcopter, as shown in Fig. 1. An Intel RealSense D435i RGB-D camera is used for both static and dynamic perception. In addition, the LiDAR Inertial Odometry (LIO) algorithm [36] is employed for accurate state estimation of the robot. The robot parameters and settings are detailed in Table I. During the experiments, we set a maximum velocity limit of 1.5 m/s along each axis to balance safety and efficiency. An occupancy map with a resolution of 0.1 m is applied to represent static obstacles, and a prediction time of 3.0 s is used for dynamic obstacle prediction.
TABLE I: Robot parameters and settings.
Max. Axis Vel. 1.5 m/s Collision Box [0.5, 0.5, 0.3]$\text{m}^{\text{3}}$
Camera FOV [86, 57]° Camera Range 5 m
Map Resolution 0.1 m Prediction Time 3.0 s
IV-A Simulation Experiments
We conduct both qualitative and quantitative experiments to evaluate the proposed navigation system. The qualitative experiments aim to demonstrate the necessity of integrating the dynamic perception and prediction modules into the system and to highlight the effectiveness of the prediction module. An example of our qualitative experiment is illustrated in Fig. 5. In this experiment, we create a specific simulation scenario where the robot is required to navigate to its goal while a dynamic obstacle moves perpendicularly across the robot’s forward path (Fig. 5a). As shown in Fig. 5b, the trajectory generation using the EGO planner [5] fails due to noisy occupancy data caused by the dynamic obstacle. In contrast, Fig. 5d demonstrates that our planner successfully generates a safe trajectory using a cleaned map and a tracked dynamic obstacle bounding box. This difference illustrates that the occupancy update in a static map is insufficiently responsive, underscoring the importance of our combined static and dynamic perception module design. When comparing the trajectories generated with and without the prediction module, Fig. 5c clearly shows that the trajectory without prediction leads to a potential future collision with the dynamic obstacle moving in the direction of the blue arrow, while the prediction module enables the generation of a trajectory that safely avoids future collisions.
@@ IV-A Simulation Experiments
The quantitative results measure the number of collisions for our system and benchmark methods across environments with different difficulty levels, as shown in Fig. 6. The white cylinders are static obstacles, and the red cylinders are dynamic obstacles that move randomly. We selected the EGO planner [5] and ViGO [9] as benchmarks. We also compared the performance of our method with and without the prediction module. Each method was tested 20 times in each environment, and the number of collisions was recorded. The results are shown in Table II. The entries represent the number of collisions, with the values in parentheses indicating the percentage relative to ViGO. Overall, our planner demonstrates the fewest collisions across all three environments. The N/A value for the EGO planner indicates that its trajectory generation fails due to noisy maps in the medium and high-density environments. Comparing the number of collisions for ViGO with our method without prediction reveals that our method achieves a lower collision rate even though ViGO uses linear prediction for dynamic obstacles. From our experimental observations, when a dynamic obstacle is close to the robot, the trajectory generated by B-spline trajectory optimization (ViGO) does not allow the robot to maneuver aggressively enough to avoid immediate collisions. In contrast, our method, which uses model predictive control, allows for more agile avoidance maneuvers. Besides, the comparison between our method with and without the prediction module shows that the proposed prediction method effectively reduces collisions, which matches the observations in the qualitative experiments.
Figure 6: Visualization of experiment environments with varying obstacle densities used to evaluate the number of collisions. Static obstacles are shown in white and dynamic obstacles are displayed in red.
TABLE II: Benchmark of the number of collisions from 20 sample runs in environments with different obstacle densities.
@@
IV-B Physical Flight Tests
To validate the performance of the system in real-world scenarios, we conducted experiments to replicate situations where the robot needs to safely operate alongside humans. We selected two indoor environments and one outdoor environment for conducting the physical flight tests, as illustrated in Fig. 7. In each environment, the robot either builds the map in real-time for straight-line navigation or uses a pre-built map with waypoints to navigate through complex structures. To simulate dynamic obstacles, we let persons intentionally walk toward the robot at a normal walking speed, blocking the robot’s paths. The left column images in Fig. 7 capture moments when the robot encounters humans, while the right column images display the generated safe trajectories for collision avoidance. Our observations confirm that the robot successfully navigates safely in all environments, demonstrating the effectiveness of the proposed method.
@@ IV-B Physical Flight Tests
Since the proposed dynamic perception module is based on a camera with a limited field of view, we also performed experiments using a motion capture system to verify obstacle avoidance ability from all directions, as shown in Fig. 8. In this experiment, the robot hovers at a fixed position while multiple pedestrians walk toward it. As illustrated in Fig. 8, the robot can successfully avoid the pedestrian and return to its original position once the pedestrian moves away.
Figure 8: Experiment on collision avoidance Using a motion capture system. The figure illustrates the robot successfully avoiding a dynamic obstacle.
@@
During these experiments, we measured the computation time for each module within the proposed navigation system, as summarized in Table III. It is important to note that all modules run in parallel on separate threads. The perception module, responsible for both static and dynamic perception, requires times of 15 ms and 27 ms, respectively. The dynamic obstacle prediction module, which includes intent and trajectory prediction, demonstrates highly efficient performance, with computation times of 0.05 ms for intent prediction and 2 ms for trajectory prediction. Finally, the planning module shows a computation time of 10 ms for generating a single MPC trajectory, and the total time for intent-based planning, including multiple MPC trajectories, is 68 ms. These results demonstrate that the system components perform within a duration suitable for real-time applications, enabling effective navigation and obstacle avoidance in dynamic environments.
TABLE III: Measurement of Proposed System Computation Time.
Proposed System Module Component Comp. Time
[…]
V Conclusion and Future Work
@@
This paper introduces a comprehensive autonomous navigation framework to enhance the safety and effectiveness of indoor UAV navigation in dynamic environments. Our approach addresses key challenges in perception and planning by incorporating a dynamic obstacle intent prediction mechanism. Specifically, we develop a perception module that efficiently detects and tracks dynamic obstacles, mitigating issues of tracking loss and occlusion during collision avoidance. We further employ a Markov Decision Process (MDP) in our intent prediction module to anticipate potential actions of dynamic obstacles and generate future trajectories. To ensure safe navigation, we propose an intent-based planning algorithm that leverages model predictive control (MPC) to compute optimal trajectories. Simulation and physical experiments validate the effectiveness of the proposed framework in enabling reliable and safe navigation of aerial robots in complex and dynamic environments. To achieve a better performance in navigating dynamic environments, our future work will focus on integrating a lightweight LiDAR to enhance the tracking range of the current system.

### v2

@@ 2 $N_{\text{pred}}\leftarrow\text{Prediction Time Steps}$;
$p^{o},v^{o},a^{o},s^{o}\leftarrow\textbf{obsInfo}(o)$ ;
  $\triangleright$ pos, vel, accel, size
@@
13                     $\mathcal{S}_{\text{control}}\leftarrow{\textbf{getLinearControlSet}}(a^{o})$
@@
15                          $\mathcal{S}_{\text{control}}\leftarrow{\textbf{getTurningControlSet}}(a^{o},\mathbb{I}^{o})$
@@
IV-C MPC Formulation for Trajectory Generation
@@
The collision constraints (Eqn. IV-C) ensure that the robot avoids collisions with static and dynamic obstacles. Static obstacles are initially represented in an occupancy voxel map, while dynamic obstacles use axis-aligned 3D bounding boxes. To unify these representations, we convert static obstacles into oriented 3D bounding boxes using a hierarchical clustering method. First, we convert the occupancy voxels into a point cloud and then apply DBSCAN clustering to determine bounding box centroids and dimensions for obstacle clusters. These bounding boxes are then refined by iteratively applying K-means++, starting with two centroids at the box corners, until the point-cloud density within each box exceeds a specified threshold. Finally, the 2D orientation angle $\phi$ of each static bounding box is discretely adjusted to maximize point-cloud density. With both static and dynamic obstacles uniformly represented as 3D bounding boxes, the collision constraint for each obstacle can be expressed as:
@@
IV-D Intent-Based Trajectory Planning
@@ 18           $\textbf{return}\ \sigma_{\text{output}}$;
The proposed algorithm (Alg. 2) generates multiple trajectories corresponding to different obstacle intent combinations and selects the optimal one based on an evaluation system. Initially, static obstacles are clustered into a set of bounding boxes $\mathcal{S}_{so}$, and the set of risky dynamic obstacles $\mathcal{S}_{do}$ near the robot is identified (Lines 2-2). Next, the top $N_{ic}$ intent combinations $S_{ic}$ for different obstacles are determined based on their current intent probability distributions. Each intent combination (labeled as $\mathcal{IC}$ in Line 2) represents a possible future action scenario for all obstacles, and the algorithm focuses on the most likely $N_{ic}$ scenarios. The likelihood of each combination is computed by multiplying the individual intent probabilities of each obstacle, assuming the distributions are independent across obstacles. For a single intent combination, the trajectory-level prediction is obtained for each obstacle (Line 2), and MPC is used to generate a collision-free trajectory $\sigma_{\text{safeTraj}}$ for navigation (Line 2). The score of the generated trajectory is calculated by combining the intent combination probability $P_{ic}$ with the raw score from the evaluation system (Lines 2-2). The algorithm selects the highest-score trajectory for execution.
@@
Both the consistency score and the detouring score are capped at a maximum allowable value to prevent extremely high scores when the denominators approach zero. However, due to the motion uncertainties of dynamic obstacles, predictions cannot always be guaranteed to be valid. To further reduce the risk of collisions while preserving consistency with previous trajectories, the trajectory is preferred when it maintains a sufficient distance from obstacles. The safety score is therefore determined by the distance to static and dynamic obstacles under the most likely intent combination:
@@
V Result and Discussion
To evaluate the proposed method, we conduct simulation and flight tests in dynamic environments. An Intel RealSense D435i camera is used for both static and dynamic perception, and the LiDAR Inertial Odometry (LIO) [31] is adopted for state estimation. Physical flight tests run on an NVIDIA Orin NX onboard computer. Experiments use a maximum velocity of 1.5m/s per axis, a 0.1m resolution occupancy map, and a 3.0s prediction time for dynamic obstacles. A time step of 0.1s and a horizon of 30 are used for MPC.
V-A Simulation Experiments
The qualitative experiments, in which dynamic obstacles follow predefined polynomial trajectories to simulate human motion with varying velocity and intent, demonstrate the need to integrate dynamic perception and prediction while highlighting the prediction module’s effectiveness. For example, Fig. 5a illustrates a scenario where the robot navigates to its goal while a dynamic obstacle crosses its path. Using the EGO planner [5], trajectory generation fails due to noisy occupancy data (Fig. 5b). In contrast, our planner successfully generates a safe trajectory with a cleaned map and tracked dynamic obstacle bounding box (Fig. 5d), showing that static occupancy updates are insufficient and emphasizing our combined perception module. When comparing the trajectories generated with and without the prediction module, Fig. 5c clearly shows that the trajectory without prediction leads to a potential future collision with the dynamic obstacle moving in the direction of the blue arrow, while the prediction module enables the generation of a trajectory that safely avoids future collisions.
@@ V-A Simulation Experiments
Figure 6: Experiment environments with varying obstacle densities. Static obstacles are shown in white, and dynamic obstacles are displayed in red.
The quantitative results show the number of collisions for our system and benchmark methods across environments with different difficulty levels (Fig. 6). The difficulty metric, survivability, are calculated as described in [32]. The white cylinders represent static obstacles, while the red cylinders denote dynamic obstacles moving at a constant velocity toward shifting goals, resulting in non-linear trajectories. We benchmark against the EGO planner [5], ViGO [9], and CC-MPC [7] and evaluate our method with and without the prediction module and safety score. Each method was tested 20 times per environment, with collision counts shown in Table I (percentages relative to ViGO in parentheses). Overall, our planner demonstrates the fewest collisions across all three environments. The N/A value for the EGO planner indicates that its trajectory generation fails due to noisy maps in the medium and high-density environments. Even without the prediction module, our approach outperforms ViGO, whose B-spline optimization lacks agility for close-proximity obstacles, whereas our MPC-based method enables more responsive avoidance. CC-MPC maintains a relatively low collision rate but experiences solver failures in high-density environments due to its strict chance constraints. Instead of relying on such strict constraints, our method improves safety by planning with multiple intents, leading to fewer solver failures in high-density environments, thus fewer collisions. The comparison of our method with and without the prediction module shows that prediction effectively reduces collisions, aligning with qualitative observations. Similarly, removing the safety score results in slightly higher collision rates in low and medium-density environments but a significant increase in high-density scenarios, highlighting its role in trajectory selection. Additionally, comparing our method without the prediction module to the version with prediction (but without the safety score) confirms that prediction significantly lowers collision rates across all environments.
TABLE I: Benchmark of the number of collisions from 20 sample runs in environments with different obstacle densities.
@@
V-B Physical Flight Tests
To validate the system’s real-world performance, we conducted physical flight tests in two indoor and one outdoor environments, as shown in Fig.7. In each environment, the robot either builds the map in real-time for straight-line navigation or uses a pre-built map with waypoints to navigate through complex structures. To simulate dynamic obstacles, we let persons intentionally walk toward the robot at a normal walking speed, blocking the robot’s paths. The left column of Fig. 7 shows the robot encountering humans, while the right column displays the safe trajectories generated for collision avoidance. Our observations confirm that the robot navigates safely in all environments, demonstrating the effectiveness of the proposed method.
@@ V-B Physical Flight Tests
Since the proposed dynamic perception module uses a camera with a limited field of view, we also performed experiments using a motion capture system to verify obstacle avoidance ability from all directions (see supplementary video). In the experiment, the robot hovers in place while multiple pedestrians approach, successfully avoiding them and returning to its original position once they move away.
@@
During experiments, we measured the computation time for each module of the proposed navigation system. All modules run in parallel on separate threads. The perception module takes 15ms for static and 27ms for dynamic perception. The dynamic obstacle prediction module requires 0.05ms for intent prediction and 2ms for trajectory prediction. The planning module computes a single MPC trajectory in 10ms. The computational time for intent-based planning, which includes multiple MPC trajectories, depends on the number of obstacles but remains manageable due to the limited number of local obstacles. On average, intent-based planning takes 68ms. These results demonstrate that the system operates within real-time constraints, enabling effective navigation and obstacle avoidance in dynamic environments.
V-C Intent-based Trajectory Prediction Evaluation
We evaluated our intent-based trajectory prediction using motion capture experiments to obtain ground truth trajectories and measure average and final displacement errors (ADE and FDE) (Table II). We let persons walk in front of the robot’s camera to simulate collision avoidance scenarios. The experiment scenarios were divided into two categories: with static obstacles and without static obstacles. For each scenario, over 100 dynamic obstacle trajectories were collected and compared between linear prediction and our intent-based method, utilizing both ground truth and perception module data. Our intent-based method generates multiple trajectories based on different intents, so we calculated displacement errors by selecting the trajectory closest to the ground truth. This evaluation approach is reasonable because our planning algorithm considers all predicted trajectories when generating plans. Therefore, if any of the predicted trajectories is accurate, the planner can use it to ensure effective navigation. Table II shows that our intent-based prediction method outperforms linear prediction by achieving lower ADE and FDE which enables more effective collision avoidance. Additionally, the prediction has a slightly higher error when using the perception module compared to the ground truth obstacle data. Furthermore, environments without static obstacles present higher prediction errors due to the less constrained motion of dynamic obstacles, which increases the difficulty of accurate trajectory prediction.
TABLE II: Evaluation of average and final displacement errors for linear and intent-based trajectory prediction methods.
Experiments w/o static obstacles w/ static obstacles
Metrics [meter] ADE FDE ADE FDE
Linear w/ mocap 1.339 3.211 0.840 1.965
Ours w/ mocap 0.699 1.264 0.543 1.064
Linear w/ perception 1.554 3.611 0.899 2.116
Ours w/ perception 0.805 1.397 0.552 1.104
VI Conclusion and Future Work
@@
This paper introduces an autonomous navigation framework that enhances the safety and effectiveness of UAV navigation in dynamic environments. Our approach addresses key challenges in perception and planning by integrating a dynamic obstacle intent prediction mechanism. Specifically, we use a perception module to efficiently detect and track dynamic obstacles and an intent-based prediction module to forecast their actions and trajectories. To reduce collisions, we propose an intent-based planning algorithm that leverages model predictive control (MPC) to compute optimal trajectories. Simulation experiments and physical flight tests confirm that the proposed framework improves the safety of navigation in dynamic environments. Our future work will focus on integrating a LiDAR to extend the perception range.